### v1

# AED-PADA:Improving Generalizability of Adversarial Example Detection via Principal Adversarial Domain Adaptation

Abstract
Adversarial example detection, which can be conveniently applied in many scenarios, is important in the area of adversarial defense. Unfortunately, existing detection methods suffer from poor generalization performance, because their training process usually relies on the examples generated from a single known adversarial attack and there exists a large discrepancy between the training and unseen testing adversarial examples. To address this issue, we propose a novel method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). Specifically, our approach identifies the Principal Adversarial Domains (PADs), i.e., a combination of features of the adversarial examples from different attacks, which possesses large coverage of the entire adversarial feature space. Then, we pioneer to exploit multi-source domain adaptation in adversarial example detection with PADs as source domains. Experiments demonstrate the superior generalization ability of our proposed AED-PADA. Note that this superiority is particularly achieved in challenging scenarios characterized by employing the minimal magnitude constraint for the perturbations.
1 Introduction
Recently, Deep Neural Networks (DNNs) have been playing prominent roles in many applications. Unfortunately, considerable studies have demonstrated that DNNs can be easily deceived if certain imperceptible perturbations are introduced to their inputs [?????]. These perturbed inputs, also known as adversarial examples, have enforced DNNs to produce erroneous decision and developed into a significant security concern in safety sensitive scenarios, such as autonomous driving [?] and medical diagnosis [?].
Figure 1: Schematic illustration of our proposed work. (a) represents the entire adversarial feature space as an instance, which contains 9 adversarial domains $\{A_{1},A_{2},\ldots,A_{9}\}$. (b) represents the mechanism of existing detection methods, which usually performs the training via a single source domain to detect the examples in the unseen adversarial domain, e.g. $T~{}(T=A_{9})$. (c) is a straightforward solution to improve generalization ability via randomly selecting multiple source domains. (d) presents the intuition of our work, which constructs PADs to possess a larger coverage of the entire feature space to create more potential overlaps with the target domain, which can better improve the generalization ability.
Nowadays, adversarial training has been proved to be an effective adversarial defense strategy [???]. However, it requires to possess enough knowledge about the classification models and necessitates substantial computational costs to retrain the classification models. On the contrary, adversarial example detection, a.k.a. adversarial detection, defends the adversarial attacks by distinguishing whether the inputs are benign or manipulated. The mechanism of this type of methods can be efficiently deploy to defend many applications without the requirement of extra knowledge about the core model.
Generalization ability is vital for adversarial detection methods in real-world scenarios, because these methods tend to encounter unseen attacks and they are desired to perform consistently. Current detection techniques [?????] typically achieve considerable generalization ability over a few conventional attacks, such as BIM [?], PGD [?] and C&W [?]. However, we empirically observe that these methods exhibit instability and inadequate performance against recent attacks, such as SSA [?], Jitter [?] and ILA-DA [?]. These methods tend to give unsatisfactory generalization performance, because their training process usually relies on a single known adversarial attack and they have not been developed from the perspective of boosting generalization ability.
In this paper, we provide a new perspective to further analyze the generalization ability of adversarial detection methods. Specifically, we assume that all the examples generated from adversarial attacks share a high dimensional manifold, and the features of the examples generated from each attack form a distinct domain, named Adversarial Domain (AD), as depicted in Fig. 1(a). Formally, Adversarial Domain (AD) of a particular adversarial attack is defined as the cumulative representations of all the adversarial examples, which are generated by that attack. Existing detection methods typically select a random single attack to generate the training samples, i.e., they only select one AD as the source domain, as shown in Fig. 1(b). Apparently, there is few intersections between the source and unseen target domains, which usually induce poor generalization performance. Besides, the randomness inherent in the selection of source domain will induce considerable fluctuations in generalization performance.
A straightforward solution to the above issues, as shown in Fig. 1(c), is to randomly select multiple ADs as the source domains. This strategy tends to create a larger overlap with the target domain(s) and thereby improves the generalization performance of adversarial detection. Nonetheless, this strategy also induces uncertainty, and the selected similar ADs bring additional training costs without enough performance gains.
To further address the above problems, we propose a novel detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). As shown in Fig. 1(d), by selecting Principal Adversarial Domains (PADs) as the source domains, which can significantly enlarges the coverage of the entire adversarial feature space and creates larger overlap with the target domain, we can offer superior generalization performance.
Specifically, AED-PADA contains two stages, i.e., Principal Adversarial Domains Identification (PADI) and Principal Adversarial Domain Adaptation (PADA). In the stage of PADI, since the discrepancies between the adversarial examples from various adversarial attacks are quite different compared to these of the ordinary classification tasks, we exploit adversarial supervised contrastive learning(Adv-SCL) to construct distinguishable ADs. Then, we construct AD clustering to identify similar ADs. To select the most representative ADs of each cluster to form PADs, we propose a Coverage of Entire Feature Space (CEFS) metric. With our CEFS metric, the formed PADs possess a board coverage of the entire feature space, and thus improve the likelihood of capturing the location of unknown target AD(s). In the stage of PADA, to effectively utilize the rich knowledge acquired by PADs to detect the unseen adversarial examples in the target domain, we propose PADA, which is compatible with various existing Multi-source Domain Adaptation (MDA) methods, by pioneering to exploit the mechanism of MDA for adversarial example detection. Since typical MDAs only focus on extracting the semantic features from the spatial domain, we propose an adversarial feature enhancement module to extracts features from both the spatial and frequency domains to construct a more comprehensive representation of adversarial examples.
Our major contributions can be summarized as follows.
  • We propose a novel adversarial example detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation, to significantly improve the generalization performance of adversarial detection.
  • We propose Principal Adversarial Domains Identification to identify the PADs, which possess a large coverage of the entire adversarial example feature space, with the help of the constructed AD clustering and proposed CEFS metric.
  • We propose Principal Adversarial Domain Adaptation for detecting adversarial examples, by exploiting adversarial feature enhancement based Multi-source Domain Adaptation (MDA), which is compatible with various existing MDA methods. To the best of our knowledge, this is the first work which exploits MDA for adversarial example detection.
2 Related Work
Adversarial example detection.
The majority of existing adversarial example detection methods rely on statistical features [??????]. They usually assume that the benign and adversarial examples originate from different distributions, and construct detectors based on the distinct statistical characteristics of these examples. Specifically, Grosse et al. [?] utilize Maximum Mean Discrepancy for the adversarial example detection. Feinman et al. [?] carry out the detection based on Kernel Density (KD) and Bayesian-Uncertainty (BU) estimation. Ma et al. [?] exploit the concept of Local Intrinsic Dimensionality (LID) to calculate the distance between the distribution of inputs and their neighbors. Lee et al. [?] utilize Guassian Discriminant Analysis (GDA) to model the difference between the benign and adversarial samples, and differentiate them based on Mahalanobis Distance (MD). Liu et al. [?] propose a steganalysis based detection method (Steg). Tian et al. [?] reveal the inconsistency in the boundary fluctuations between the adversarial and benign examples, and construct Sensitivity Inconsistency Detector (SID) to identify the adversarial examples. Wang et al. [?] embed hidden-layer feature maps of DNNs into word vectors, and detecting adversarial examples via Sentiment Analysis (SA).
Figure 2: Our AED-PADA framework contains two stages: (a) Principal Adversarial Domains Identification, which consists of Adversarial Domain Acquisition, Adversarial Domain Clustering and Principal Adversarial Domains Selection, and (b) Principal Adversarial Domain Adaptation.
3 Methodology
To improve the generalization ability of adversarial example detection, we propose a novel adversarial example detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). The entire framework of our AED-PADA is shown in Fig. 2. AED-PADA contains two stages, Principal Adversarial Domains Identification (PADI) and Principal Adversarial Domain Adaptation (PADA). In the stage of PADI, we firstly incorporate adversarial supervised contrastive learning(Adv-SCL) to acquire distinguishable ADs. Then, we construct AD clustering to group ADs into different clusters. By proposing the Coverage of Entire Feature Space (CEFS) metric, we select the most representative ADs from each cluster to form PADs. In the stage of PADA, we propose an adversarial feature enhancement based MDA method to effectively leverage PADs to achieve adversarial detection. Note that Secs. 3.2, 3.3, and 3.4 introduce PADI while Sec. 3.5 presents PADA.
3.1 Notations
Suppose we have a labeled $C-$class classification dataset $\mathbb{D}=\{(x,y)\}_{i=1}^{N}$ with $N$ samples, where the label $y\in\{1,2,\ldots,C\}$. A classifier $f$ is trained on $\mathbb{D}$ to classify an input sample into one of the $C$ classes, $f(x)\rightarrow\mathbb{Z}_{C}$. We have a set of adversarial attacks, $\mathbb{A}=\{Adv_{1},Adv_{2},\ldots,Adv_{M}\}$, which consists of $M$ distinct adversarial attack methods. Each of them aims to deceive $f$ into assigning incorrect labels via the generated adversarial examples. $\mathbb{D}_{m}=\{(x^{\prime},y^{\prime},\psi)\}_{i=1}^{N}$ is the dataset generated by attack $Adv_{m}$, where $x^{\prime}$ and $y^{\prime}$ denote the adversarial examples and their corresponding predictions, respectively, and $\psi$ presents the attack employed to generate $x^{\prime}$. $\psi=m$, for all $\psi$ in $\mathbb{D}_{m}$. Consequently, we construct a set of adversarial examples $\mathbb{O}$, which comprises $M$ types of adversarial examples, $\mathbb{O}=\{\mathbb{D}_{1},\mathbb{D}_{2},\ldots,\mathbb{D}_{M}\}$.
3.2 Adversarial Domain Acquisition
Typically, if we extract the features of the adversarial examples, which are generated from different untargeted attacks, via common CNNs, the features tend to spread in an indistinguishable manner in the feature space. To acquire distinguishable representations of adversarial examples from different attacks, we exploit adversarial supervised contrastive learning (Adv-SCL) to extract features. Then, we form the Adversarial Domains (ADs), each of which is defined as the cumulative representations of all the adversarial examples generated from a particular adversarial attack.
Based on the supervised contrastive learning (SCL) method [?], Adv-SCL neglects the classification result $y^{\prime}$ of adversarial examples and only focuses on identifying their generation methods. Specifically, each adversarial example in $\mathbb{O}$ is characterized by its generation method, denoted as $\psi$, which serves as a key discriminant between the positive and negative samples. Here, a pair of examples from identical attack are considered as positive, while those from different attacks are considered as negative. This learning strategy amplifies the dissimilarities across examples from various attacks and generates a more appropriate representation.
As shown in Fig. 2(a), the input of AD acquisition is the adversarial example set $\mathbb{O}$. Adv-SCL consists of Encoder Network ($Enc$) and Projection Network ($Proj$). $Enc(\cdot)$ extracts a feature vector from the input image $x^{\prime}$, and $Proj(\cdot)$ further project this representation vector to an auxiliary vector $z=Proj(Enc(x^{\prime}))$, which will be discarded after training.
For $n$ randomly sampled example-label pairs, $\{x_{k}^{\prime},y_{k}^{\prime},\psi_{k}\}_{k=1}^{n}$, in $\mathbb{O}$, the corresponding batch employed for training consists of $2n$ pairs, $\left\{\widetilde{x_{l}^{\prime}},\widetilde{y_{l}^{\prime}},\widetilde{\psi_{l}}\right\}_{l=1}^{2n}$, where $\widetilde{x^{\prime}}_{2k}$ and $\widetilde{x^{\prime}}_{2k-1}$ are two views of $x^{\prime}_{k}$, and $\widetilde{\psi}_{2k-1}=\widetilde{\psi}_{2k}=\psi_{k}$. The loss of Adv-SCL is defined as,
$\mathcal{L}^{sup}=\sum_{i\in I}\mathcal{L}_{i}^{sup},$ (1)
$\mathcal{L}_{i}^{sup}=\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp(z_{i}\cdot z_{p}/\tau)}{\sum_{a\in A(i)}\exp(z_{i}\cdot z_{a}/\tau)}.$ (2)
Here, $i\in I=\{1,\ldots,2N\}$ denotes the index of the augmented samples. $A(i)$ is a subset of $I$ containing all the indexes except $i$. $P(i)\equiv\{p\in A(i):\widetilde{\psi}_{p}=\widetilde{\psi}_{i}\}$ represents the indices of positive samples in the batch except $i$, and $|P(i)|$ stands for its cardinality. $\cdot$ denotes the dot product. $\tau$ is the temperature parameter which scales the similarity values.
3.3 Adversarial Domain Clustering
After the acquisition of ADs, we obtain $\Gamma_{i}=Enc(D_{i})$ as AD of $Adv_{i}$. For $\mathbb{O}$, we construct a set of ADs, =$\mathbb{R}=\{\Gamma_{1},\Gamma_{2},\ldots,\Gamma_{M}\}$. Since different ADs tend to distribute differently in the feature space and many of them possess various portions of overlaps, it is quite difficult to directly select the most representative ADs from scratch. Then, it is vital to explore the similarities among different ADs.
To address this issue, we intend to perform the selection via two steps, i.e., clustering to assess the similarities of ADs and selecting the most representative ADs from each cluster. Then, we construct a viable solution named Adversarial Domain Clustering, as shown in Fig. 2(a). This strategy groups ADs into different clusters, by ensuring that the similarity among ADs within the same cluster is maximized, while the similarity among ADs across different clusters is minimized.
Since the samples to be clustered here are collections of features, rather than individual data points, traditional clustering methods such as K-Means [?] cannot be directly applied. Since spectral clustering [?] only requires the similarity matrix among samples, it is utilized to construct the clustering step in our AD clustering.
For the estimation of the similarity matrix $W\in R^{M\times M}$ in spectral clustering, which represents the similarities between different ADs, we propose Adversarial Domain Similarity Measurement (ADSM) based on Jensen-Shannon divergence (JSD) [?], which quantifies the similarity between two probability distributions. To compute the similarities, we transform each $\Gamma_{i}$ in $\mathbb{R}$ to $\overline{\Gamma_{i}}$ by converting $\Gamma_{i}$ into probabilities and performing normalizations. Since a smaller value of JSD between two ADs implies a smaller discrepancy in their probability distributions, $\text{ADSM}(\Gamma_{i},\Gamma_{j})$ can be computed via
$\text{ADSM}(\Gamma_{i},\Gamma_{j})=\frac{1}{\text{JSD}(\overline{\Gamma_{i}},\overline{\Gamma_{j}})}.$ (3)
By letting the element $W_{i,j}$ at the $i$-th row and $j$-th column refer to the similarity between $\Gamma_{i}$ and $\Gamma_{j}$, $W_{i,j}$ can be calculated by
$W_{i,j}=\left\{\begin{array}[]{cc}\text{ADSM}(\Gamma_{i},\Gamma_{j})&i\neq j\\ 0&i=j\end{array}.\right.$ (4)
Since the spectral clustering cannot automatically determine the optimal number of clusters, to evaluate the clustering performance and estimate the optimal number of clusters, the Calinski-Harabasz score (CH score) [?], which requires no knowledge of the cluster shape, is utilized. Note that it measures both the within-cluster and between-cluster distances, thereby offering a more comprehensive view of the clustering performance. With the help of the CH score, our AD clustering can automatically group the ADs into optimal number of clusters.
3.4 Principal Adversarial Domains Selection
After similar ADs are clustered, then the selection step can be performed. To form the most effective Principal Adversarial Domains (PADs), it is vital to select proper ADs from different clusters. Thus, we propose the Coverage of Entire Feature Space metric (CEFS) to guide the PADs selection process.
CEFS is a ratio-based metric which contains two aspects, Intra-Domain Dispersion (IDD) and Discrepancy between Adversarial Domains (DAD). IDD represents the dispersion among features within each AD, where a higher value indicates a larger coverage of the feature space. For any AD, $\Gamma_{I}=\{{\Gamma_{I}}_{1},{\Gamma_{I}}_{2},\ldots,{\Gamma_{I}}_{v}\}\in R^{v\times d}$, where $v$ and $d$ denote the number and dimension of features, respectively, IDD can be computed by
$\text{IDD}(x)=\frac{1}{v}\sum_{i=1}^{v}\left\|\frac{x_{i}}{\|x_{i}\|_{2}}-\frac{1}{v}\sum_{j=1}^{v}\frac{x_{j}}{\|x_{j}\|_{2}}\right\|_{2}.$ (5)
DAD represents the discrepancies between the two selected ADs, this discrepancy can be quantified using a distance metric $Dist$, such as Kullback-Leibler divergence [?] or Maximum Mean Discrepancy (MMD) [?]. A lower DAD value indicates a larger similarity of the two selected ADs. DAD can be calculated as,
$\text{DAD}(\Gamma_{i},\Gamma_{j})=Dist(\overline{\Gamma_{i}},\overline{\Gamma_{j}}).$ (6)
Then, CEFS can be obtained via
$\text{CEFS}=\frac{\sum_{i}^{K}\text{IDD}(\Gamma_{i})}{\text{DAD}((\Gamma_{1}\parallel\cdots\parallel\Gamma_{K}),(\Gamma_{1}\parallel\cdots\parallel\Gamma_{M}))},$ (7)
where $K$ and $M$ is the number of ADs in PADs and the number of ADs in the AD set $\mathbb{R}$ respectively, and $\parallel$ is a concatenation operation. A higher CEFS value represents a larger coverage of the entire feature space. We utilize CEFS to select PADs, which can give a larger coverage of the feature space with the same number of ADs. PADs actually enhance the probability of capturing the location of unseen ADs, thereby improving the generalization performance.
3.5 Principal Adversarial Domain Adaptation
To transfer the learned knowledge from PADs to the target domain, i.e., the adversarial examples generated from unseen methods, we propose Principal Adversarial Domain Adaptation (PADA) to detect adversarial examples, as depicted in Fig. 2(b).
Ideally, the inputs of PADA comprise source data and unseen target data. The source data contains the adversarial examples from PADs. Since the unseen target data is unavailable, we can only utilize $\mathcal{O}$, the entire set of adversarial examples, as a proxy.
PADA consists of 3 sequential components, feature extraction, feature alignment and classification. PADA is compatible with various widely used multi-source domain adaptation (MDA) methods [???].
Unfortunately, existing MDA methods only extract spatial features. [??] indicate that adversarial perturbations are more likely to be concealed in the high-frequency information of images. To capture more comprehensive features of adversarial examples, we propose an adversarial feature enhancement (AFE) module to be the feature extraction component. AFE contains both the spatial feature extraction $\mathcal{S}$ and frequency feature extraction $\mathcal{F}$ branches. For frequency feature extraction, we design a plug-and-play high-pass filters (HPF) operation to enhance the high-frequency signals and remove low-frequency components of inputs. More details of HPF can be found in the supplementary materials. Both $\mathcal{S}$ and $\mathcal{F}$ employ $Enc$ from Sec. 3.2. Then, the enhanced adversarial features, denoted as $r(x)=[\mathcal{S}(x),\mathcal{F}(x)]$, are fed into the future module.
4 Experiments
4.1 Experimental Setups
DNN backbones and datasets.
We evaluate the performance of the proposed method, by employing two widely used DNN architectures, i.e., ResNet-18  [?] and VGG-16 [?], on two popular datasets CIFAR-10 [?] and SVHN [?], by following [?].
Baseline attack methods.
To evaluate the generalization capabilities of the detection methods, it is important to consider a diverse set of attack methods, including both the earlier and recent techniques. Here, 10 earlier attack methods are utilized to generate adversarial examples for training, including FGSM [?], BIM [?], C&W [?], DeepFool [?], PGD [?], MIFGSM [?], DIM[?], ILA [?], YA-ILA [?] and SINIFGSM [?]. 7 SOTA attack methods are employed to generate adversarial examples for testing, including APGD [?], ILA-DA [?], Jitter [?], SSA [?], TIFGSM [?], VMIFGSM [?] and VNIFGSM [?].
Implementation details.
In this paper, adversarial examples are generated to carry out the untargeted white-box attack with the $l_{\infty}$ norm constraint. The maximum magnitude of adversarial perturbation is set to 2. Step size and number of iterations for adversarial attacks are set to 1/255 and 10, respectively. During PADA stage, we utilize MFSAN [?] as the basic MDA method, and employ the same training strategy as it. All the adversarial example detection methods are trained consistently for 100 epochs. To evaluate the performance of our proposed method, the widely used Accuracy is employed as the metric for the adversarial example detection task. The experiments are conducted on an NVIDIA GeForce RTX 3080Ti GPU.
More details of the experimental setups can be found in the supplementary materials.
4.2 Performance Evaluations
Here, we compare our method with 5 SOTA detection methods, including LID [?], MD [?], Steg [?], SID [?] and SA [?].
Performances of defending unseen attacks.
Table 1 presents the cross-attack detection results on 7 unseen state-of-the-art testing adversarial attacks. These results provide compelling evidence that our approach obviously outperforms the SOTA methods in terms of the generalization ability. Note that the performance of our approach (91.219%) surpasses that of MD-PADs (52.266%) by a significant margin, i.e., 38.953%, under the settings of SVHN and VGG-16. Besides, PADs may not be suitable to be directly applied to the existing methods. As can be observed, if the SOTA detection methods directly adopt PADs for training (denoted as X-PADs), their performances may not always increase. This observation further verifies the effectiveness of our PADA in our AED-PADA framework.
Dataset (Backbone) Detector Accuracy on Unseen SOTA Adversarial Attacks (%) Averaged Accuracy (%)
APGD ILA-DA Jitter SSA TIFGSM VMIFGSM VNIFGSM
CIFAR-10 (ResNet-18) LID 90.044 94.121 78.240 59.839 90.863 91.056 86.496 84.380
LID-PADs 78.903 86.068 70.045 57.330 82.278 82.333 79.978 76.705
MD 65.976 96.657 61.764 54.225 69.282 69.653 69.726 69.612
MD-PADs 63.371 97.970 62.058 51.501 68.100 68.814 69.351 68.738
Steg 83.521 94.692 86.759 58.064 90.496 90.814 92.057 85.200
Steg-PADs 84.085 94.155 88.885 62.305 90.715 90.860 91.690 86.099
SID 86.113 60.999 74.629 53.732 87.323 87.330 82.100 76.032
SID-PADs 86.520 61.820 74.860 54.010 88.285 88.220 82.935 76.664
SA 84.907 90.731 87.365 81.323 89.275 90.013 91.967 87.940
SA-PADs 89.125 93.485 91.925 88.885 94.76 94.875 95.595 92.664
AED-PADA (Ours) 90.545 97.855 97.065 84.255 97.700 97.765 97.675 94.694
CIFAR-10 (VGG-16) LID 85.024 93.357 75.092 58.639 88.201 88.173 81.540 81.432
LID-PADs 70.493 89.403 65.190 58.588 75.553 75.408 73.500 72.590
MD 51.322 72.013 50.904 52.012 51.809 51.799 50.963 54.403
MD-PADs 50.033 81.779 50.028 51.349 50.005 50.856 51.349 55.057
Steg 79.280 92.756 88.466 57.470 88.854 88.891 89.390 83.586
Steg-PADs 79.900 92.265 88.705 59.360 89.015 88.800 89.235 83.897
SID 83.235 66.047 73.260 53.373 84.743 84.582 76.210 74.493
SID-PADs 83.105 66.240 73.230 53.375 84.605 84.385 76.145 74.441
SA 80.198 85.036 81.003 63.984 85.356 86.293 87.627 81.357
SA-PADs 79.720 83.825 77.985 63.335 83.765 84.475 85.755 79.873
AED-PADA (Ours) 84.370 94.365 93.260 65.765 93.135 93.410 93.545 88.264
SVHN (ResNet-18) LID 66.431 87.940 63.697 60.324 71.403 71.384 64.952 69.447
LID-PADs 61.849 86.846 60.551 58.515 69.581 69.639 64.182 67.309
MD 61.167 66.439 59.906 58.270 67.625 67.609 60.870 63.127
MD-PADs 60.768 50.190 59.637 57.967 67.257 67.337 60.688 60.549
Steg 71.591 51.509 94.268 59.865 96.943 96.874 95.954 81.000
Steg-PADs 73.460 50.860 97.115 74.360 98.005 98.040 97.575 84.202
SID 68.859 92.207 63.763 56.571 74.074 73.995 64.150 70.517
SID-PADs 69.670 93.420 64.155 56.800 75.285 75.120 64.780 71.319
SA 68.092 73.066 65.288 69.276 82.758 82.823 82.467 74.824
SA-PADs 69.490 73.100 63.610 66.550 78.155 78.220 77.785 72.416
AED-PADA (Ours) 74.455 99.730 99.730 87.305 99.730 99.730 99.730 94.344
SVHN (VGG-16) LID 65.645 96.605 62.731 64.252 74.205 74.263 70.296 72.571
LID-PADs 69.710 97.154 63.626 63.955 76.258 76.306 70.119 73.875
MD 51.511 58.333 51.233 51.859 51.870 52.268 51.465 52.648
MD-PADs 51.920 56.484 51.578 52.745 51.103 50.285 51.748 52.266
Steg 70.586 98.661 93.882 58.496 95.774 95.575 94.549 86.789
Steg-PADs 72.925 98.975 97.700 58.905 98.485 98.480 98.085 89.079
SID 68.159 69.813 61.320 60.722 74.468 74.454 67.213 68.021
SID-PADs 68.365 62.450 61.145 56.790 70.075 70.080 63.470 64.625
SA 65.862 72.558 63.304 67.745 72.087 72.183 73.811 69.650
SA-PADs 65.060 69.765 59.365 62.605 66.795 66.955 67.385 65.419
AED-PADA (Ours) 73.500 99.220 98.995 68.990 99.295 99.310 99.220 91.219
Table 1: Comparison of the generalization performances between the state-of-the-art adversarial detection methods and our AED-PADA. For the methods without the ‘-PADs’ mark, since they are trained on single attack, each of their results is the averaged detection accuracy with 10 training attacks being utilized one after another in training. The methods with the ‘-PADs’ mark are trained by the adversarial examples in PADs, which are consistent with our AED-PADA.
Generalization performances across different backbones and datasets.
To thoroughly evaluate the generalization ability of adversarial detection methods, it is also vital to assess their performances in the scenarios where the training and testing adversarial examples are from different backbones and datasets. A detection method with better cross-attack, cross-backbone, and cross-dataset capabilities tends to be better suited for practical applications in complex environments. Table 2 displays the generalization performances of different detection methods across different backbones and datasets. In general, results from both scenarios demonstrate that our method exhibits the best generalization ability. Since the structures tend to vary significantly among different types of backbones, which will result in different feature dimensions between training and testing, both LID and MD are not applicable in cross-backbone experiment, i.e., they lack the capability to generalize across backbones, because they utilize the intermediate features of the backbones for detection.
Detector Averaged Accuracy (%)
Backbone Dataset
[…]
LID - - 71.619 80.233
MD - - 62.437 61.429
Steg 82.894 85.157 53.249 78.155
SID 74.169 76.458 68.928 73.442
SA 80.751 87.773 58.910 51.694
Ours 89.301 91.624 82.964 85.113
Table 2: Comparison of generalization performances across different backbones and datasets. The cross-backbone results are obtained on CIFAR-10, and the cross-dataset results are obtained when the backbone is ResNet-18.
Origin Detector Averaged Accuracy (%) CEFS($\uparrow$)
Same Cluster FGSM+SINIFGSM+PGD 90.444 -
BIM+ILA+YA-ILA 92.513 -
MIFGSM+SINIFGSM+DIM 93.494 -
Cross Cluster DIM+BIM+C&W 94.694 80.856
MIFGSM+BIM+C&W 94.621 72.625
SINIFGSM+BIM+C&W 94.624 58.649
FGSM+BIM+C&W 94.600 58.543
PGD+BIM+C&W 94.222 50.669
FGSM+BIM+DeepFool 93.407 50.142
FGSM+ILA+DeepFool 92.539 49.818
PGD+ILA+DeepFool 92.429 48.015
Table 3: Comparison of generalization performances when different source domains are employed in the training process. The dataset is CIFAR-10, and the backbone is ResNet-18. The result of AD clustering is {FGSM, PGD, DIM, MIFGSM, SINIFGSM}-{BIM, ILA, YA-ILA}-{C&W, DeepFool}. ‘Same Cluster’ and ‘Cross Cluster’ respectively represent that the ADs employed as the sources domains are selected from the same AD cluster and different AD clusters.
Dataset (Backbone) Averaged Accuracy (%)
$K=2$ $K=3$ $K=4$ $K=5$ $K=6$ $K=7$
CIFAR-10 (ResNet-18) 91.936 94.694 94.784 94.571 94.561 95.018
CIFAR-10 (VGG-16) 87.946 88.264 87.964 87.991 88.104 88.498
SVHN (ResNet-18) 91.497 94.344 91.257 91.726 94.999 94.801
SVHN (VGG-16) 89.183 89.269 91.219 90.056 90.026 90.713
Table 4: Results with different number of clusters $K$. The underlined and bolded values respectively represent the results of our AED-PADA, in which the number of clusters is automatically determined based on the CH score, and the best averaged accuracies.
Dataset Feature Extraction Averaged Accuracy (%) Average
ResNet-18 VGG-16
CIFAR-10 spatial 93.411 88.251 90.831
freq 93.379 86.755 90.067
spatial + freq 94.694 88.264 91.479
SVHN spatial 94.612 89.721 92.166
freq 88.157 91.580 89.869
spatial + freq 94.344 91.219 92.781
Table 5: Results of different feature extractions in PADA of our AED-PADA.
Dataset Backbone Averaged Accuracy (%)
MFSAN (Ours) M3DA-$\beta$
CIFAR-10 ResNet-18 94.694 94.976
VGG-16 88.864 87.819
SVHN ResNet-18 94.344 93.892
VGG-16 91.219 90.960
Table 6: Results of our framework with different MDA methods.
4.3 Analysis
Effectiveness of AD clustering and PADs Selection.
As shown in Table 3, we set up two groups of experiments, i.e., selecting ADs from Same Cluster and Cross Cluster, to verify the effectiveness of our AD clustering and PADs Selection, against the 7 unseen attacks. According to Table 3, for ‘Same Cluster’, the best and worst averaged results are 90.444% and 93.494%, respectively. For ‘Cross Cluster’, the best, worst, and averaged values are 94.694%, 92.429%, and 93.892%, respectively. Then, we can obtain three observations: (1) The averaged results of ‘Cross Cluster’ selection is higher than the best result of ‘Same Cluster’ selection, and the best result of ‘Cross Cluster’ selection is also superior. This verifies the effectiveness of our AD Clustering. Apparently, source domains from ‘Cross Cluster’ can certainly enhance the generalization ability of the detection methods. (2) Although ‘Cross Cluster’ selection in general achieves better results, it cannot guarantee that the randomly selected ADs (from ‘Cross Cluster’s) always give better performance than the results of ‘Same Cluster’ selected ADs. When the PADs with a very low CEFS score are employed as the source domains for detection, the generalization performance may be worse than that of ‘Same Cluster’ selection. This observation further verifies the effectiveness of our PADs Selection. (3) For the majority of results, a higher CEFS value of PADs induces a higher averaged accuracy of the detection method across different target domains. This observation suggests that the proposed CEFS is a proper guide for selecting PADs, i.e., a higher CEFS score indicates that the corresponding PADs can cover a larger proportion of the entire feature space, and thus gives a better performance.
The effect of different numbers of clusters.
Here, we apply AD Clustering to 10 adversarial attack methods in the training set and present the impact of different cluster numbers $K$ (the number of adversarial attacks in PADs). As depicted in Table 4, the detection performance increase when $K$ increases, until a certain value of $K$ is achieved. This trend is attributed to the utilization of an increased number of ADs as the source domains, which undeniably provides larger coverage of the entire feature space. Note that we can employ the CH score to automatically determine the number of clusters, which is included in AD Clustering of our AED-PADA, and these results are underlined in Table 4. As can be observed, our automatic CH score based method can yield superior performance on SVHN with VGG-16 being the backbone, and give comparable performance to the best result on other settings. Although the best result with manually selecting $K$ can exceeds our automatic method for 0.2%-0.3% in terms of accuracy, on average, it requires 2.3×2.3$\times$2.3 × more parameters and gives a 3×3$\times$3 × slower speed. This suggests that our method actually achieves a better balance between the training costs and generalization performances.
Effectiveness of the adversarial feature enhancement in PADA.
In this experiment, we utilize 3 distinct feature extractions, i.e., spatial feature extraction, frequency feature extraction, and the adversarial feature enhancement (spatial + freq) in our framework. Still 7 unseen attacks are employed and the averaged results are reported in Table 5. As can be observed, our adversarial feature enhancement (spatial + freq) exhibit superior performances than both the spatial feature extraction and frequency feature extraction, which verifies the effectiveness of our work.
The compatibility with different MDA methods.
Here, we employ two widely used MDA methods, MFSAN [?] and M3DA-$\beta$ [?], to demonstrate the compatibility of our framework with the existing MDA methods. Table 6 presents that the performances of employing both the MFSAN and M3DA-$\beta$ can surpass the existing SOTA adversarial detection methods. This verifies that our framework possesses excellent compatibilities with existing MDA methods. Besides, since MFSAN outperforms M3DA-$\beta$ in the majority of experiments, we select MFSAN as the default MDA component in our PADA.
5 Conclusion
In this paper, we proposed a novel and effective adversarial example detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation, to improve the generalization ability of adversarial detection. Specifically, AED-PADA contains two stages, i.e., Principal Adversarial Domains Identification and Principal Adversarial Domain Adaptation. In PADI, we acquired ADs from scratch and constructed PADs as the source domains for PADA. In PADA, we constructed an adversarial feature enhancement based Multi-source Domain Adaptation, which is compatible with various existing MDA methods, to effectively leverage PADs to achieve adversarial detection. Experimental results proved the superiority of our work, compared to the state-of-the-art detection methods.